OpenStudy (anonymous):

65% of what number is 39? A. 25.35 B. 60 C. 166.7 D. 2535

OpenStudy (onepieceftw):

Really easy once you know how, just multply 39 by the decimal form of the percent. 65% is 0.65. What's 39 times 0.65?

OpenStudy (anonymous):

\(x\times \frac{69}{100}=39\) x=?

OpenStudy (anonymous):

x=26/91?